OpenStudy (anonymous):

use the given information to find the exact value find sin(A-B) using the sine difference identity cos A= 1/3,0

OpenStudy (anonymous):

help me plz

OpenStudy (anonymous):

help me plz

OpenStudy (anonymous):

Trivial. Just use the sin(A-B) = sinAcosB - sinBCosA formula. You can work out sin(A) from cos(A) (and vice versa) using cos^2(A) + sin^2(A) = 1, which implies sin(A) = +/- sqrt(1-cos^2(A). Use bounds on where A and B are to decide whether each is the positive or negative root.

OpenStudy (anonymous):

i really dont get this plz explain me more in a board

OpenStudy (anonymous):

No, sorry. I gave you everything you nee.d Now do your own homework. <3

OpenStudy (anonymous):

plz help me ranga sir plz

OpenStudy (ranga):

sin(A-B) = sinAcosB - sinBCosA cosA = 1/3 sinB = -1/2 sin(A-B) = sinAcosB - sinBCosA = sin(A-B) = sinAcosB - (-1/2)(1/3) = sinAcosB + 1/6  Find sinB and cosA from the diagram and substitute.

OpenStudy (anonymous):

how do i substitute sir

OpenStudy (ranga):

First find the length of the missing side in the two triangles marked with a question mark. That is, What is PQ in the first triangle and SB in the second triangle?

OpenStudy (anonymous):

i dont no how to do that

OpenStudy (ranga):

You should get the basics first before attempting this problem. The figure shows right triangles. AQ^2 + PQ^2 = AP^2 1^2 + PQ^2 = 3^2 Find PQ. (This is the same as using sin^2A + cos^2A = 1. But I showed you the triangle so you can know what sine and cosine represents in a triangle).

OpenStudy (anonymous):

2^3

OpenStudy (ranga):

1^2 + PQ^2 = 3^2 1 + PQ^2 = 9 PQ^2 = 9 - 1 = 8 PQ = sqrt(8) = sqrt(4*2) = 2sqrt(2) sin(A) = opposite / hypotenuse = PQ / AP = 2sqrt(2) / 3. This is what you will substitute for sin(A). Use the second triangle to find cos(B) in a similar way.

OpenStudy (anonymous):

hmm i got for the pother one sqyrt (2)/2

OpenStudy (anonymous):

is it right

OpenStudy (anonymous):

hello sir

OpenStudy (anonymous):

sqrt(3)/2

OpenStudy (ranga):

yes. And since B is in the fourth quadrant cos(B) is positive. sin(A-B) = sinAcosB + 1/6 sin(A) = 2sqrt(2) / 3 cos(B) = sqrt(3)/2 Substitute into sin(A-B) = sinAcosB + 1/6

OpenStudy (anonymous):

so my answer is 2sqrt(2)/3 and sqrt(3)/2 right sir

OpenStudy (anonymous):